The power dissipated in a resistor is 240 watts when a voltage of 220 volts is across it. If the voltage is cut in half, what is the power dissipation in the resistor?


1
Expert's answer
2021-12-27T08:21:49-0500

"P= \frac{V^2}{R}"


Where P= power

V= voltage

R= resistance


For a constant resistance, "P\propto V^2"

When voltage is cut in "\frac{1}{2}"

Power is cut in "(\frac{1}{2})^2"


New Power= "(\frac{1}{2})^2×240=60W"